Sup suppose the level of autonomous investment in an economy is 2,000 shillings and the consumption function is c=800 +0.25Y find the equilibrium levels



What will be the increase in national income if the investment increases by 250


Suppose that the level of planned investment is to the ceilings and the saving function is given by S=-800 + 0.25Y ,find the equilibrium levels of income


Suppose the consumption function is given by C=200+0.2Y ,find the equilibrium levels of income


With is a graph discuss the ranges of the aggregate supply curve

1
Expert's answer
2021-11-24T12:31:17-0500

As per the guidelines we are allowed to answer the first three subparts only. Please post the remaining questions separately so that those can also be answered. Thank You.

Income(Y) = Consumption + Investment  ........................1

Income(Y) = Consumption + Savings  ....................2

From 1 and 2, we get:

Investment = Savings

Step 2

autonomous investment = 2,000 shillings

consumption function , c=800 + 0.25Y

Y = 800 + 0.25Y + 2000

⇒Y - 0.25Y = 2800

⇒Y =1

0.75


10.75×

× 2800  (where 1.75 is multiplier and 2800 is autonomous expenditure)

⇒Y = 3733.33

Income=3733.33shillings

Consumption=800 + 0.25(3733.33) = 1733.33shillings

 

Step 3

When investment increases by 250, national income increases by:

multiplier×

×250=1

0.75


10.75×250

×250 = 333.33shillings

Step 4

When the level of planned investment is to the ceilings and the saving function is S=-800 + 0.25Y, then

Y = Consumption + Savings

⇒Y = 1733.33 + (-800 + 0.25Y)

⇒Y - 0.25Y = 933.33

⇒Y = 1244.44

Income = 1244.44shillings